write three terms of general solution
{pi/6+kpi}

Answer by Theo(13342)

your general solution is:
y = (pi/6) + (k * pi)
it's easier to do the math if the denominators are the same.
set pi = 6pi/6 and you can add much easier.
you will get:
when k = 0, pi/6 = pi/6 + 0*6pi/6 = pi/6 + 0 = pi/6.
when k = 1, pi/6 + 1*6pi/6 = pi/6 + 6pi/6 = 7pi/6
when k = 2, pi/6 + 2*6pi/6 = pi/6 + 12pi/6 = 13pi/6
when k = 3, pi/6 + 3*6pi/6 = pi/6 + 18pi/6 = 19pi/6